# Cooperative Sheaf Neural Networks

### v2

@@ Abstract
Sheaf neural networks (SNNs) leverage cellular sheaves to induce flexible diffusion processes on graphs, generalizing the diffusion mechanism of classical graph neural networks. While SNNs have been shown to cope well with heterophilic tasks and alleviate oversmoothing, we show that there is further room for improving sheaf diffusion. More specifically, we argue that SNNs do not allow nodes to independently choose how they cooperate with their neighbors, i.e., whether they convey and/or gather information to/from their neighbors. To address this issue, we first introduce the notion of cellular sheaves over directed graphs and characterize their in- and out-degree Laplacians. We then leverage our construction to propose Cooperative Sheaf Neural Network (CSNN). Additionally, we formally characterize its receptive field and prove that it allows nodes to selectively attend (listen) to arbitrarily far nodes while ignoring all others in their path, which is key to alleviating oversquashing. Our results on synthetic data empirically substantiate our claims, showing that CSNN can handle long-range interactions without succumbing to oversquashing. We also show that CSNN performs strongly in heterophilic node classification and long-range graph classification benchmarks.
@@ 1 Introduction
Our theoretical results show that CSNN allows for nodes to selectively listen to other arbitrarily distant nodes, which is a desirable trait to alleviate over-squashing. Our results on synthetic data specifically designed to induce over-squashing (Alon and Yahav, 2021) substantiate our claims, showcasing CSNNs’ superior potential to handle long-range dependencies. In addition, extensive real-world experiments on 11 node-classification benchmarks and 2 long-range graph-classification tasks demonstrate that CSNNs consistently outperform existing SNNs and cooperative GNNs.
@@
    We carry an extensive experimental campaign to validate the effectiveness of CSNNs, encompassing both synthetic and real-world tasks. Our experiments on synthetic data show that CSNNs is remarkably capable of mitigating over-squashing and modeling long-range dependencies. Meanwhile, results on over 13 real-world tasks show CSNNs consistently surpass prior sheaf-based models and cooperative GNNs.
@@ 2 Background
That being said, note that $\ker L_{{\mathcal{F}}}=\{\mathbf{X}\in C^{0}(G,{\mathcal{F}})\>|\>\mathcal{F}_{i\unlhd e}\mathbf{x}_{i}=\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}\,\forall e=\{i,j\}\in E(G)\}$. This can be understood as the space of public agreement between all pairs of neighboring nodes $i$ and $j$. Note that $i$ and $j$ can have distinct opinions about the same topic on their respective private opinion spaces $\mathcal{F}(i)$ and $\mathcal{F}(j)$; however, when they publicly discuss this topic, they may prefer to not manifest their true opinion. This apparent consensus lies in $\operatorname{Ker}L_{{\mathcal{F}}}$.
@@ 3 Cellular sheaves for directed graphs
We note that if $(G,{\mathcal{F}})$ is the trivial sheaf, then $L^{\text{out}}_{\mathcal{F}}=(L^{\text{out}})^{\top}$ and $L^{\text{in}}_{\mathcal{F}}=L^{\text{in}}$. Flat vector bundles over directed graphs. We can also improve the parameter efficiency of cellular sheaves over directed graphs using flat vector bundles. Since the graphs are directed, we need to distinguish between edges with identical endpoints but with different orientations. Thus, for each node $i$, we assign a source conformal map ${\mathbf{S}}_{i}$ and a target conformal map ${\mathbf{T}}_{i}$, and set $\mathcal{F}_{i\unlhd ij}={\mathbf{S}}_{i}$ and $\mathcal{F}_{i\unlhd ji}={\mathbf{T}}_{i}$ for all neighbors $j$ of $i$.
@@ 4 Cooperative Sheaf Neural Networks
Note these matrices have a block structure, with diagonals $({L^{\text{in}}_{\mathcal{F}}})^{\top}_{ii}=\sum{\mathbf{T}}_{i}^{\top}{\mathbf{T}}_{i}$, $({L^{\text{out}}_{\mathcal{F}}})_{ii}=\sum{\mathbf{S}}_{i}^{\top}{\mathbf{S}}_{i}$, and remaining blocks $(L^{\text{in}}_{{\mathcal{F}}})^{\top}_{ij}=(L^{\text{out}}_{{\mathcal{F}}})_{ij}=-{\mathbf{T}}^{\top}_{i}{\mathbf{S}}_{j}$. We point out that, since conformal maps are of the form ${\mathbf{S}}_{i}=C_{{\mathbf{S}}_{i}}Q_{i}$ and ${\mathbf{T}}_{i}=C_{{\mathbf{T}}_{i}}R_{i}$, for some orthogonal matrices $Q_{i}$, $R_{i}$ and scalars $C_{{\mathbf{S}}_{i}},\>C_{{\mathbf{T}}_{i}}$, the block diagonals are just scalars times an identity matrix, hence the normalization is both numerically stable and computationally efficient.
@@ 4.1 Analysis
Thus, to show a model achieves cooperative behavior, we must ensure the following: (a) if $i$ does not listen, then its update cannot depend on $\mathbf{x}_{j}$, $\forall j\in V,j\neq i$; and (b) if $i$ has a non-propagating neighbor $k$, then its update cannot depend on $\mathbf{x}_{k}$. Proposition˜4.1 shows CSNN satisfies these conditions.
@@
In each layer $t$, the features of a node can be affected by features of nodes up to $2t$-hop neighbors.
@@
Observe that the derivative of $\mathbf{x}_{1}^{(3)}$ in relation to $\mathbf{x}_{4}^{(0)}$ can be as high as the values of the non-zero ${\mathbf{T}}_{i}$ and ${\mathbf{S}}_{i}$ permit. This shows our model can mitigate over-squashing, which refers to the failure of an information propagating to distance nodes. Di Giovanni et al. (2023) and Topping et al. (2022), studied over-squashing in message passing neural networks through a bound on the Jacobian $\left|\nicefrac{{\partial\mathbf{x}_{i}^{(t)}}}{{\partial\mathbf{x}_{j}^{(0)}}}\right|\leq c^{t}\hat{A}^{t}_{ij},$ where $t$ is the layer, $c$ is a constant that depends on the architecture of the model, and $\hat{A}$ the normalized adjacency matrix. Moreover, over-squashing occurs when we have a small derivative $\nicefrac{{\partial\mathbf{x}_{i}^{(t)}}}{{\partial\mathbf{x}_{j}^{(0)}}}$, since it means that after $t$ layers, the feature at $i$ is mostly insensitive to the information initially contained at $j$, i.e., the information was not propagated properly. Proposition˜4.3 states that the feature at node $i$ can be sensitive to the information initially contained at node $j$, independently of the distance, given enough layers and an appropriate configuration of restriction maps.
@@ 5 Related works
Cooperative GNNs. Finkelshtein et al. (2024) were the first to propose flexibilizing message passing by allowing nodes to choose how they cooperate with each other. Each layer of their model, CO-GNN, employs an additional GNN that chooses the cooperation the action each node takes. While CO-GNNs can be employed with arbitrary base and action networks, their main caveat is that training can become increasingly difficult as these networks become more complex — the grid of hyper-parameters grows considerably and the stochastic nature of the action network may affect model selection. Different from CO-GNN, our CSNN does not rely on discrete actions and can smoothly modulate between cooperative behavior patterns.
Sheaf Neural Networks. Besides works on SNNs for graph data with real-valued node features, recent works have also expanded the literature to deal with heterogenous edge types  (Braithwaite et al., 2024), hypergraphs (Duta et al., 2023), and node features living on Riemann manifolds  (Battiloro et al., 2024). Additionally, while recent works on SNNs learn restriction maps in an end-to-end fashion, there are also prior works in which they are manually constructed  (Hansen and Ghrist, 2019) or computed as a pre-processing step (Barbero et al., 2022).
Quiver Laplacians. Sumray et al. (2024) defined sheaf Laplacians over quivers. In their context, quivers are directed graphs with self-loops. The in- and out-Laplacians we defined here are not particular cases of the Laplacians over quivers: they obtain a positive semi-definite matrix while our Laplacians can have complex eigenvalues with negative real parts.
@@ 6 Experiments
We provide both synthetic and real-world experiments to evaluate the performance of CSNN, including node- and graph-level prediction tasks. Section˜6.1 assess CSNN’s capacity to circumvent over-squashing using the NeighborsMatch benchmark proposed by Alon and Yahav (2021). Section˜6.2 presents experiments on eleven node classification tasks, showcasing the effectiveness of CSNN for heterophilic graphs. Finally, Section˜6.3 consider the Peptides datasets from the Long Range Graph Benchmark (Dwivedi et al., 2022) to substantiate the capability of our model to mitigate under-reaching and over-squashing on real-world graph classification.
@@ 6.1 Over-squashing
In order to verify our theoretical results on the capacity of CSNN to alleviate over-squashing, we reproduce the NeighborsMatch problem proposed by Alon and Yahav (2021), using the same framework. The datasets consist of binary trees of depth $r$, with the root node as the target, the leaves containing its possible labels, and the leaf with the same number of neighbors as the target node containing its true label. We provide the parameters used for this task in Appendix˜B.
@@
Figure˜4 shows GCN (Kipf and Welling, 2017) and GIN (Xu et al., 2019) fail to fit the datasets starting from $r=4$ and GAT (Veličković et al., 2018) and GGNN (Li et al., 2016) fail to fit the datasets starting from $r=5$. These models suffer from over-squashing and are not able to distinguish between different training examples, while the CSNN model reaches perfect accuracy for all tested $r$.
@@ 6.2 Node Classification
Experimental setting. As baselines for benchmarks in Platonov et al. (2023), we use GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018) and GT (Shi et al., 2021), together with the variations GAT-sep and GT-sep, which concatenate the representation of a node to the mean of its neighbors instead of summing them (Zhu et al., 2020). These are all classical baselines used in Platonov et al. (2023) to compare against GNN architectures specifically developed for heterophilic settings, and that achieve the best performance in most cases. We also compare CSNN against recent models such as CO-GNN, NSD, and BuNN (Bamberger et al., 2025). Results for BuNN on Squirrel and Chameleon are not available, since we do not have access to their code.
For the remaining datasets, we compare against the classical GCN, GAT and SAGE; the models specifically tailored for heterophilic data GGCN (Yan et al., 2022), Geom-GCN (Pei et al., 2020), H2GCN (Zhu et al., 2020), GPRGNN (Chien et al., 2020), FAGCN (Bo et al., 2021) and MixHop (Abu-El-Haija et al., 2019); and the models GCNII (Chen et al., 2020) and PairNorm (Zhao and Akoglu, 2020) designed to alleviate oversmoothing. We use the 10 fixed splits proposed by Platonov et al. (2023) and Pei et al. (2020). We refer to Appendix˜B for further implementation details.
Results. Table˜1 and Table˜2 show that CSNN is the best-performing method in 9 out of 11 datasets. These results highlight our model’s capacity to deal with heterophilic graphs of different sizes and heterophily levels. We note CSNN often outperforms both NSD and CO-GNN by a significant margin.
@@ 8 Ethics and Reproducibility Statements
Reproducibility Statement. Aiming to secure reproducibility of our work, we provide proofs of our theoretical results and in experiment detail in Appendix˜A and Appendix˜B. Moreover, we will provide a public code once the review process is complete.
@@ Appendix A Proofs
If $\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}=0$ for every $j$, then $L_{\mathcal{F}}(\mathbf{X})_{j}=0$.
A.2 Proof of Proposition 4.1
@@
Appendix B Experiment details

### v3

@@ Abstract
Sheaf neural networks (SNNs) leverage cellular sheaves to induce flexible diffusion processes on graphs, generalizing the diffusion mechanism of classical graph neural networks. While SNNs have been shown to cope well with heterophilic tasks and alleviate oversmoothing, we show that there is further room for improving sheaf diffusion. More specifically, we argue that SNNs do not allow nodes to independently choose how they cooperate with their neighbors, i.e., whether they convey and/or gather information to/from their neighbors. To address this issue, we first introduce the notion of cellular sheaves over directed graphs and characterize their in- and out-degree Laplacians. We then leverage our construction to propose Cooperative Sheaf Neural Network (CSNN). Additionally, we formally characterize its receptive field and prove that it allows nodes to selectively attend (listen) to arbitrarily far nodes while ignoring all others in their path, which is key to alleviating oversquashing. Our results on synthetic data empirically substantiate our claims, showing that CSNN can handle long-range interactions while avoiding oversquashing. We also show that CSNN performs strongly in heterophilic node classification and long-range graph classification benchmarks.
@@ 1 Introduction
Our theoretical results show that CSNN allows for nodes to selectively listen to other arbitrarily distant nodes, which is a desirable trait to alleviate over-squashing. Our results on synthetic data specifically designed to induce over-squashing (Alon and Yahav, 2021) substantiate our claims, showcasing CSNNs’ superior potential to handle long-range dependencies. In addition, extensive real-world experiments on 11 node-classification benchmarks and 2 long-range graph-classification tasks demonstrate that CSNN typically outperforms existing SNNs and cooperative GNNs.
@@
    We carry an extensive experimental campaign to validate the effectiveness of CSNNs, encompassing both synthetic and real-world tasks. Our experiments on synthetic data show that CSNNs is remarkably capable of mitigating over-squashing and modeling long-range dependencies. Meanwhile, results on over 13 real-world tasks show CSNNs typically outperform prior sheaf-based models and cooperative GNNs.
@@ 2 Background
That being said, note that $\ker L_{{\mathcal{F}}}=\{\mathbf{X}\in C^{0}(G,{\mathcal{F}})\>|\>\mathcal{F}_{i\unlhd e}\mathbf{x}_{i}=\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}\,\forall e=\{i,j\}\in E(G)\}$. This can be understood as the space of public agreement between all pairs of neighboring nodes $i$ and $j$. Note that $i$ and $j$ can have distinct opinions about the same topic on their respective private opinion spaces $\mathcal{F}(i)$ and $\mathcal{F}(j)$; however, when they publicly discuss this topic, they may prefer to not manifest their true opinion. Alternatively, since the edge stalks may be different from the node stalks, some topics of the private opinion spaces do not need to be discussed at all. In both cases, the apparent consensus lies in $\operatorname{Ker}L_{{\mathcal{F}}}$.
@@ 3 Cellular sheaves for directed graphs
We note that if $(G,{\mathcal{F}})$ is the trivial sheaf, then $L^{\text{out}}_{\mathcal{F}}=(L^{\text{out}})^{\top}$ and $L^{\text{in}}_{\mathcal{F}}=L^{\text{in}}$.
Flat vector bundles over directed graphs. We can also improve the parameter efficiency of cellular sheaves over directed graphs using flat vector bundles. Since the graphs are directed, we need to distinguish between edges with identical endpoints but with different orientations. Thus, for each node $i$, we assign a source conformal map ${\mathbf{S}}_{i}$ and a target conformal map ${\mathbf{T}}_{i}$, and set $\mathcal{F}_{i\unlhd ij}={\mathbf{S}}_{i}$ and $\mathcal{F}_{i\unlhd ji}={\mathbf{T}}_{i}$ for all neighbors $j$ of $i$.
@@ 4 Cooperative Sheaf Neural Networks
Note these matrices have a block structure, with diagonals $({L^{\text{in}}_{\mathcal{F}}})^{\top}_{ii}=\sum{\mathbf{T}}_{i}^{\top}{\mathbf{T}}_{i}$, $({L^{\text{out}}_{\mathcal{F}}})_{ii}=\sum{\mathbf{S}}_{i}^{\top}{\mathbf{S}}_{i}$, and remaining blocks $(L^{\text{in}}_{{\mathcal{F}}})^{\top}_{ij}=(L^{\text{out}}_{{\mathcal{F}}})_{ij}=-{\mathbf{T}}^{\top}_{i}{\mathbf{S}}_{j}$. We point out that, since conformal maps are of the form ${\mathbf{S}}_{i}=C_{{\mathbf{S}}_{i}}Q_{i}$ and ${\mathbf{T}}_{i}=C_{{\mathbf{T}}_{i}}R_{i}$, for some orthogonal matrices $Q_{i}$, $R_{i}$ and scalars $C_{{\mathbf{S}}_{i}},\>C_{{\mathbf{T}}_{i}}$, computing their inverses and normalizing the Laplacian becomes trivial. In this case, block scaling simplifies to a block matrix of scalars time identity and the normalization is both numerically stable and computationally efficient.
@@ 4.1 Analysis
Thus, to show a model achieves cooperative behavior, we must ensure the following: (a) if $i$ does not listen, then its update cannot depend on $\mathbf{x}_{j}$, $\forall j\in V,j\neq i$; and (b) if $i$ has a non-propagating neighbor $k$, then its update cannot depend on $\mathbf{x}_{k}$. Proposition˜4.1 shows that CSNN satisfies these conditions, while Figure˜3 illustrates them and delineates limitations of the NSD model.
Figure 3: Given a graph $(a)$ we illustrate the consequences of preventing node 33 from listening. For NSD $(b)$, this means $L_{\mathcal{F}}(\mathbf{X})_{3}$ must not depend on $\mathbf{x_{j}},$ for $j=1,2,4$ implying $\mathcal{F}_{j\unlhd e}=0$ and leading to $L_{\mathcal{F}}(\mathbf{X})_{j}$ not depending on $\mathbf{x_{3}}$, preventing node 33 from propagating information. In CSNN $(c)$, we can set ${\mathbf{T}}_{3}=0$. Provided ${\mathbf{S}}_{3}\neq 0$, outbound communication is possible.
@@
In each layer $t$, the features of a node can be affected by the features of nodes up to $2t$-hops.
@@
This configuration shows that although CSNN can achieve $2$-hop neighbors, it can also refrain from this behavior to only access $1$-hop neighbors per layer. Moreover, this flexibility indicates there are multiple forms to stablish communication between two distant nodes while ignoring others.
@@
Observe that the derivative of $\mathbf{x}_{1}^{(3)}$ in relation to $\mathbf{x}_{4}^{(0)}$ can be as high as the values of the non-zero ${\mathbf{T}}_{i}$ and ${\mathbf{S}}_{i}$ permit. This shows our model can mitigate over-squashing, which refers to the failure of an information propagating to distance nodes. Di Giovanni et al. (2023) and Topping et al. (2022), studied over-squashing in message passing neural networks through a bound on the Jacobian $\left|\nicefrac{{\partial\mathbf{x}_{i}^{(t)}}}{{\partial\mathbf{x}_{j}^{(0)}}}\right|\leq c^{t}\hat{A}^{t}_{ij},$ where $t$ is the layer, $c$ is a constant that depends on the architecture of the model, and $\hat{A}$ the normalized adjacency matrix. Moreover, over-squashing occurs when we have a small derivative $\nicefrac{{\partial\mathbf{x}_{i}^{(t)}}}{{\partial\mathbf{x}_{j}^{(0)}}}$, since it means that after $t$ layers, the feature at $i$ is mostly insensitive to the information initially contained at $j$, i.e., the information was not propagated properly. Proposition˜4.3 states that the feature at node $i$ can be sensitive to the information initially contained at node $j$, independently of the distance, given enough layers and an appropriate configuration of restriction maps. Consequently, Proposition˜4.3 suggests that the restriction maps regulating the sensibility between distant nodes can provide higher upper bounds to $\nicefrac{{\partial\mathbf{x}_{i}^{(t)}}}{{\partial\mathbf{x}_{j}^{(0)}}}$ while decreasing the value of $\nicefrac{{\partial\mathbf{x}_{i}^{(t)}}}{{\partial\mathbf{x}_{k}^{(0)}}}$ for other nodes $k$.
@@ 5 Related works
Cooperative GNNs. Finkelshtein et al. (2024) were the first to propose flexibilizing message passing by allowing nodes to choose how they cooperate with each other. Each layer of their model, CO-GNN, employs an additional GNN that chooses an action for each node. While CO-GNNs can be employed with arbitrary base and action networks, their main caveat is that training can become increasingly difficult as these networks become more complex — the grid of hyper-parameters grows considerably and the stochastic nature of the action network may affect model selection. Different from CO-GNN, our CSNN does not rely on discrete actions and can smoothly modulate between cooperative behavior patterns.
Sheaf Neural Networks. Besides works on SNNs for graph data with real-valued node features, recent works have expanded the literature to accommodate heterogeneous edge types  (Braithwaite et al., 2024), hypergraphs (Duta et al., 2023), nonlinear Sheaf Laplacians  (Zaghen et al., 2024), and node features living on Riemann manifolds  (Battiloro et al., 2024). While recent works on SNNs learn restriction maps in an end-to-end fashion, there are also prior works in which they are manually constructed  (Hansen and Ghrist, 2019) or computed as a pre-processing step (Barbero et al., 2022).
Quiver Laplacians. Sumray et al. (2024) propose sheaf Laplacians over quivers (directed graphs w/ self-loops) to improve feature selection on tabular data, with no learning component involved. The in- and out-Laplacians we defined here are not particular cases of the Laplacians over quivers. The former are positive semi-definite matrices, while our Laplacians may have complex eigenvalues with negative real parts.
@@ 6 Experiments
We provide both synthetic and real-world experiments to evaluate the performance of CSNN, including node- and graph-level prediction tasks. Section˜6.1 assess CSNN’s capacity to circumvent over-squashing using the NeighborsMatch benchmark proposed by Alon and Yahav (2021). Section˜6.2 presents experiments on eleven node classification tasks, showcasing the effectiveness of CSNN for heterophilic graphs. Finally, Section˜6.3 consider the Peptides datasets from the Long Range Graph Benchmark (Dwivedi et al., 2022) to substantiate the capability of our model to mitigate under-reaching and over-squashing on real-world graph classification. We also provide additional experiments in Appendix˜B.
@@ 6.1 Over-squashing
In order to verify our theoretical results on the capacity of CSNN to alleviate over-squashing, we reproduce the NeighborsMatch problem proposed by Alon and Yahav (2021), using the same framework. The datasets consist of binary trees of depth $r$, with the root node as the target, the leaves containing its possible labels, and the leaf with the same number of neighbors as the target node containing its true label. We provide the parameters used for this task in Appendix˜C.
Figure˜5 shows GCN (Kipf and Welling, 2017) and GIN (Xu et al., 2019) fail to fit the datasets starting from $r=4$ and GAT (Veličković et al., 2018) and GGNN (Li et al., 2016) fail to fit the datasets starting from $r=5$. These models suffer from over-squashing and are not able to distinguish between different training examples, while the CSNN model reaches perfect accuracy for all tested $r$.
@@
Comparison against other sheaf models. Notably, CSNN outperforms other sheaf methods in this task. BuNN reports $100\%$ accuracy until $r=6$. Then it drops to $71\%$ and $42\%$ for $r=7$ and $r=8$, respectively, as reported in Bamberger et al. (2025). For NSD with orthogonal maps, we obtained $100\%$ accuracy when $r=2$, $91\%$ for $r=3$, and then a sharp drop to $5\%$ when $r=4$.
@@ 6.2 Node Classification
Experimental setting. As baselines for benchmarks in Platonov et al. (2023), we use GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018) and GT (Shi et al., 2021), together with the variations GAT-sep and GT-sep, which concatenate the representation of a node to the mean of its neighbors instead of summing them (Zhu et al., 2020). These are all classical baselines used in Platonov et al. (2023) to compare against GNN architectures specifically developed for heterophilic settings, and that achieve the best performance in most cases. We also compare CSNN against recent models such as CO-GNN, NSD, and BuNN (Bamberger et al., 2025). Results for BuNN on Squirrel, Chameleon, and the datasets in Table 2 are not available, since we do not have access to their code.
For the remaining datasets, we compare against the classical GCN, GAT and SAGE; the models specifically tailored for heterophilic data GGCN (Yan et al., 2022), Geom-GCN (Pei et al., 2020), H2GCN (Zhu et al., 2020), GPRGNN (Chien et al., 2020), FAGCN (Bo et al., 2021), FSGNN (Maurya et al., 2022), GloGNN Li et al. (2022), ACMGCN (Luan et al., 2022), and MixHop (Abu-El-Haija et al., 2019); and the models GCNII (Chen et al., 2020) and PairNorm (Zhao and Akoglu, 2020) designed to alleviate oversmoothing. We use the 10 fixed splits proposed by Platonov et al. (2023) and Pei et al. (2020). We refer to Appendix˜C for further implementation details.
Results. Table˜1 and Table˜2 show that CSNN is the best-performing method in 9 out of 11 datasets. These results highlight our model’s capacity to deal with heterophilic graphs of different sizes and heterophily levels. We note CSNN often outperforms both NSD and CO-GNN in Table˜1. While we report the results in Table˜2 for completeness, we note they exhibit high variance — in accordance to the findings of Platonov et al. (2023), which highlight that the small scale of these datasets may incur unstable and statistically insignificant results.
@@ 8 Ethics and Reproducibility Statements
Reproducibility Statement. Aiming to secure reproducibility of our work, we provide proofs of our theoretical results and in experiment detail in Appendix˜A and Appendix˜C. Moreover, we will provide a public code once the review process is complete.
@@ Appendix A Proofs
If $\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}=0$ for every $j$, then $L_{\mathcal{F}}(\mathbf{X})_{j}=0$. ∎
A.2 Remark regarding Proposition 3.1 and nonlinear Sheaf Laplacian
If the sheaf Laplacian is nonlinear as in Zaghen et al. [2024], i.e, $L_{\mathcal{F}}(\mathbf{X})_{i}=\sum_{i,j\unlhd e}\mathcal{F}_{i\unlhd e}^{\top}\phi_{e}\left(\mathcal{F}_{i\unlhd e}\mathbf{x}_{i}-\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}\right)$, where $\phi_{e}:{\mathcal{F}}(e)\to{\mathcal{F}}(e)$ is a continuous function for each edge $e$, then saying that $L_{\mathcal{F}}(\mathbf{X})_{i}$ does not depend of $\mathbf{x}_{j}$ means $\mathcal{F}_{i\unlhd e}^{\top}\phi_{e}\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}=0$. Then the proof holds similarly to the above.
A.3 Proof of Proposition 4.1
@@
Appendix B Additional Experiments
In this section, we provide other two experiments. The first is a real-world large dataset and the second is a noisy implementation of Example˜4.4, to illustrate how Proposition˜4.3 may work in practice.
Table 4: Results on the Penn94 dataset.
Models Penn94
Edge Homophily 0.47
MLP 73.61 $\pm$ 0.40
GCN 82.47 $\pm$ 0.27
GAT 81.53 $\pm$ 0.55
MixHop 83.47 $\pm$ 0.71
GCNII 82.92 $\pm$ 0.59
H2GCN 81.31 $\pm$ 0.60
WRGAT 74.32 $\pm$ 0.53
GPR-GNN 81.38 $\pm$ 0.16
ACM-GCN 82.52 $\pm$ 0.96
LINKX 84.71 $\pm$ 0.52
GloGNN 85.57 $\pm$ 0.35
CSNN 86.00 $\pm$ 0.39
Real-world. We run CSNN on the Penn94, a dataset with $41.554$ nodes, $1.362.229$ edges, and whose feature dimension is $5$, against multiple baselines as reported in Li et al. [2022] to illustrate CSNN’s performance in a larger ambiguous heterophilic [Luan et al., 2024] dataset. We highlight that CSNN also outperform baselines on Squirrel, another dataset classified as ambiguous heterophilic.
Synthetic. In the following, we consider a path graph with four nodes as in Example˜4.4 where the features are initialized in a similar fashion of the TreeNeighborsMatch task, but we consider that all nodes have a number of “blue neighbors" to add noisy information along the path, while the source and target (nodes 3 and 0) have the same number of these neighbors. The goal is to transfer information from node 3 to node 0.
Figure˜6 shows an actual run of CSNN in such graph: if the norm of ${\mathbf{T}}_{i}$ is close to zero, then the arrow pointing to $i$ is not draw. Analogously to ${\mathbf{S}}_{i}$. If both are close to zero, the node is isolated, but if none of them are we consider the following:
  • If ${\mathbf{T}}_{i}{\mathbf{S}}_{j}$ (in the in-Laplacian) and ${\mathbf{T}}_{i}{\mathbf{S}}_{j}$ (in the out-Laplacian) are both close to 0, then the edge $(j,i)$ does not exist
  • If ${\mathbf{T}}_{i}{\mathbf{S}}_{j}$ (in the out-Laplacian) and ${\mathbf{T}}_{i}^{\top}{\mathbf{T}}_{i}$ are both close to 0, then edge $(j,i)$ does not exist.
We observe restriction maps going to zero in a different configuration than the one exhibited in Example˜4.4 but still sending $x_{3}^{(0)}$ to node 0 in the second layer, while suppressing other node features.
Figure 6: Illustration of running CSNN in a path graph with four nodes along three layers.
Technically, the significance of a node $i$ over $j$ in a given layer $t$ is measured by the norm of the the term that multiplies $x_{i}^{(t)}$ in the expression of $((\Delta_{{\mathcal{F}}(t)}^{in})^{\top}\Delta_{{\mathcal{F}}(t)}^{out})_{j}$. For instance, in Figure˜6, the restriction maps provided that $1$ listens to 0 and no one else, so it can be affected only by $x_{1}^{(0)}$ and by $x_{0}^{(0)}$. A calculation shows that $x_{1}^{(0)}$ is multiplied by a quantity about $0.7$ while $x_{0}^{(0)}$ is multiplied by $132$, approximately. This guarantees that in layer $2$, when the message goes from node $1$ to 0, the feature $x_{1}^{(0)}$ is not relevant and, in practice, is ignored.
Appendix C Experiment details